OpenStudy (zasharra):

flvs students!!! please help. (03.05 Module Three Quiz.)

OpenStudy (zasharra):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (zasharra):

@Ghostgate

OpenStudy (zasharra):

for f(x) i think i got one

OpenStudy (ghostgate):

Alright, what do you think it is? Also, I don't know how to necessarily write these out, I could probably think of something, but it would take awhile for me.

OpenStudy (zasharra):

that's fine i really need help! is mine right (i really dont know this stuff)

OpenStudy (ghostgate):

If I recall correctly, f(x), g(x), h(x), and j(x) are just stand alone numbers. I did some research real quick on what a quadratic equation is, a quadratic equation deals with squares, which was shown above, so... for instance it would look something like: 3x^2 + 10 = 342 - This is just an example. So... now to solve your problem.

OpenStudy (ghostgate):

Well before I go any further, does that help you out?

OpenStudy (zasharra):

what do you mean stand alone numbers?

OpenStudy (ghostgate):

Well, actually they are apart of a graph, in other words, it means that: f(x) = 2x^2 + 5x = 4y - What this means is that you would use this equation for a graph, if I recall correctly, so you wouldn't do anything with 'f(x)'. Also, if I may ask, what grade level of Math are you doing?

OpenStudy (zasharra):

10th

OpenStudy (ghostgate):

Oh, I'm doing 12th, which is Consumer Math, because I didn't want to take Trigonometry or Calculus or whatever alternative I would have had. Lol

OpenStudy (ghostgate):

Alright, so does this help or do you want me to assist you in your problem? In other words, do you think you can do the rest or not?

OpenStudy (zasharra):

so what would be my final answer? f(x) = g(x) = h(x) = j(x) = (The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.)

OpenStudy (ghostgate):

I would think it would be something like this: f(x) = 2x^2 - 5x^2 - Because a difference of squares should be subtracting two squares. g(x) = x^2 + 4x^2 - Because the sum of course is adding. That's for the first two, I can't entirely figure out what the others mean by 'square trinomial' and 'GCF' out of it when I don't know how to find that out with square variables.

OpenStudy (zasharra):

@Ghostgate did you ever figure out the other two?

OpenStudy (zasharra):

@amistre64